OpenStudy (anonymous):

What is the quotient written in scientific notation? 2 * 10^21 _________ 8 * 10^3 A. 0.25 * 10^7 B. 0.25 * 10^18 C. 2.5 * 10^6 D. 2.5 * 10^17

OpenStudy (loser66):

the first one is 2/8 = 1/4 = 0.25 , just let 10^ (something) aside and calculate the front part . then calculate the 10^ (something) I mean